OpenStudy (anonymous):

use sum-to-product to find EXACT VALUE of sin150degrees + sin30degrees

OpenStudy (anonymous):

why can't you just add them?

OpenStudy (anonymous):

\[\sin(150)=\frac{1}{2}\] \[\sin(30)=\frac{1}{2}\] \[\frac{1}{2}+\frac{1}{2}=1\] finis

OpenStudy (anonymous):

It seemed too easy lol. and I kept getting either 1 or 0 .

OpenStudy (anonymous):

it is pretty darned easy considering the other ones

OpenStudy (anonymous):

Well it says I have to use the sum- to- product formula

OpenStudy (anonymous):

then pretend to use it and write 1 as the answer

OpenStudy (anonymous):

xD kay . I got a verifying identity i can't solve >.<

OpenStudy (anonymous):

ok ok so we have to have \[\sin(\alpha+\beta)+\sin(\alpha-\beta)=2\sin(\alpha)\cos(\alpha)\] right?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

so lets see \[\alpha +\beta=150\] \[\alpha -\beta=30\] solve and get \[2\alpha =180\implies \alpha =90\] and so \[\beta =60\]

OpenStudy (anonymous):

and in it goes: \[2\sin(90)\cos(60)=2\times 1\times \frac{1}{2}=1\]

OpenStudy (anonymous):

lotta work for that simple answer i would say

OpenStudy (anonymous):

xD I appreciate it though